Exercise 11-3 Make or Buy Decision (LO11-3) Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment.Exercise 11-3 Make or Buy Decision (LO11-3] Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment.Exercise 11-3 Make or Buy Decision (LO11-3] Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment.Exercise 11-3 Make or Buy Decision (LO11-3] Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment.
Answer #1

(1) Relevant cost of making = DM + DL + Variable mfr OH + Supervisor salary

= $12 + $8 + $2 + $3 = $25

Relevant cost of 12000 carburetors = $25 * 12000 = $300000

Cost of buying = $30

Cost of 12000 carburetors = $30 * 12000 = $360000

Financial (disadvantage) = $300000 - $3600000 = ($60000)

(2) No, outside suppliers offer should not be accepted.

(3) Segment margin of new product = $120000

Financial advantage = $120000 + ($60000) = $60000

(4) yes, outside suppliers offer should not be accepted.